# Should you use LLMs to simulate opinions? Quality checks for early-stage deliberation

### v3

@@ Discussion
Applying our quality checks, we uncovered significant logical inconsistencies: specifically, most models (80%) produced “average” opinions more extreme than demographic-specific predictions, violating basic statistical logic. Unlike logical inconsistencies of LLMs exposed in previous research—for example, mutually incompatible factual beliefs (Kassner et al. 2021), self‑inconsistent chain‑of‑thought traces (Wang et al. 2022), violations of basic propositional logic (Ghosh et al. 2025), or within‑passage contradictions (Mündler et al. 2024)—these errors indicate fundamental logical flaws at the distributional level. Practitioners cannot simply acknowledge and proceed despite these errors; they must reconsider or even abandon flawed simulation strategies entirely.
Further, our checks reveal systematic discrepancies with stakeholder expectations. Older models (like GPT-3.5-turbo) tend to exaggerate opinion differences even on obviously non-divisive claims, risking harmful stereotyping or inefficient resource allocation. Conversely, newer models (GPT-4, LLaMa3, Mistral, Titan) often obscure genuine demographic divides on controversial topics, potentially masking critical areas affecting particular groups. Importantly, we find that the magnitude of gender-level differences observed in QC2a correlates with the direction of the errors observed in QC2b, suggesting that testing clear-cut cases aligned with stakeholder priors—even in the absence of extensive human data—provides a valuable preliminary quality check of LLM reliability.
@@
Organizations should prioritize testing whether model outputs align with clear common-sense or domain-specific expectations, paying close attention to discrepancies with their justified priors. For instance, an error of exaggeration of differences between groups on a non-divisive issue is likely a more serious error than an error of exaggeration on an issue known to be divisive. Moreover, while our current metrics focus on mean differences in opinion distributions, some stakeholders may prioritize other distributional properties—such as variance or skew of opinions (Bui et al. 2025; Wang, Morgenstern, and Dickerson 2025). Our framework is flexible: the test statistics we propose can be adapted to target other distributional moments.
@@ Conclusion
Appendix 1: Literature Review Table
@@
Out of One, Many (Argyle et al. 2023) Political Attitudes Prompt Engineering: Backstories ANES GPT-3 Positive. Useful for simulating opinion distributions.
Donald Trumps (Jiang, Wei, and Zhang 2024) Voting Prompt Engineering: Backstories + structural ANES, WVS ChatGPT Positive. Backstory prompts produce outputs that match voting patterns.
Synthesizing Public Opinion (Karanjai et al. 2025) Political Attitudes In-Context Learning: Knowledge Injection CES Gemma2, LLaMa3 Positive. Incorporating knowledge in prompts improves over backstory.
Mixture of Personas (Bui et al. 2025) Sentiment Prompt Engineering: Probabilistic Mixture of Personas IMDB, Yelp, AGNews LLaMa3 Positive. Better alignment and within-group diversity than backstory prompting.
Modeling Targeted Populations (Namikoshi et al. 2024) EV attitudes Fine-tuning: ID & Demographics BEV Survey LLaMa2 Positive. Fine-tuning improved KL-divergence and RMSE with subgroups.
Aligning with Whom? (Sun et al. 2023) Politeness Prompt Engineering: Backstories POPQUORN GPT-3.5, GPT-4 Mixed. Aligns more with White participants than Black/Asian ones.
Whose Opinions? (Santurkar et al. 2023) Political Attitudes Prompt Engineering: Backstories OpinionQA GPT-3.5 Mixed. Matches younger/liberal demos; some steering helps.
Climate Opinion (Lee et al. 2024) Environmental attitudes Prompt Engineering: Backstories + covariates Novel. Nat’l sample GPT-3.5, GPT-4 Mixed. Demographics alone insufficient; psych vars improve results.
Synthetic Survey Data (Bisbee et al. 2024) Feeling thermometers Prompt Engineering: Backstories ANES GPT-3.5-Turbo Mixed. Similar means, reduced variance; prompt/time sensitive.
Beyond Demographics (Orlikowski et al. 2025) Content moderation Fine-tuning: IDs & Demographics MINT, POPQUORN, DICES LLaMa 3 Mixed. Models learn individual opinions, but fail at demographic-level.
Chameleons? (Mingmeng, He, and Trotta 2024) Political Attitudes Prompt Engineering: Backstories ESS GPT-3.5, Llama 2/3, Mistral, Deepseek Negative. Highly sensitive to prompt; misrepresents group-level opinions.
LLMs Flatten Identities (Wang, Morgenstern, and Dickerson 2025) Political Attitudes Prompt Engineering: Backstories Novel (n=3200) GPT-4, Llama 2, WizardLM Negative. Misrepresents minorities; flattens within-group variation.
Cross-Cultural Simulation (Qu and Wang 2024) Political Attitudes Prompt Engineering: Backstories WVS GPT-3.5-Turbo Negative. Performs better on U.S. samples; shows liberal bias.
Appendix 2a: Data Collection Procedure - TopicMisinfo Dataset
Using Amazon SageMaker Ground Truth, we hired American workers and compensated them at a rate of \$15-20 per hour to annotate various claims. We asked them to annotate each claim according to different important axes of prioritization. Specifically, we asked for annotations of a claim’s along four dimensions: (1) “Prioritization”: need for prioritization in fact-checking (based on their own subjective criteria); (2) “General Public”: interest to the general public; (3) “Group Harm”: potential to harm certain demographic groups; and (4) “Perceived Falsehood”: perceived degree of veracity. Table Appendix 2a: Data Collection Procedure - TopicMisinfo Dataset shows the questions that annotators were asked. If workers identified a claim as harmful to a specific group(s), we asked them to identify which group(s) may be harmed. After workers completed the task, we asked them to fill out a demographic survey.
@@
Appendix 2b: Example Claims from Dataset
@@
Appendix 3: Descriptive Analysis of Dataset
@@
Appendix 4: Prompts for Experiments
@@
Appendix 5a: Pseudocode for Quality Check 1a.
@@
Appendix 5b: Pseudocode for Quality Check 1b.
@@
8:    Initialise abs_diffs $\leftarrow[\;]$
@@
20: Initialise bootstrap_results $\leftarrow[\;]$
@@
Appendix 5c: Replication of Quality Check 1 including Non-binary Gender
@@
Appendix 6a: Fine-tuning approach for benchmark evaluation
@@
Appendix 6b: In-context learning approach for benchmark evaluation
@@
Appendix 7: LLM Positioning of Opinions
This chart complements the results from Experiment 3, as seen in Tables 1 and 3, by showing the positioning of LLM-generated opinions on a Likert scale for men and women.
@@
Appendix 8: Salient Examples of Opinion Differences Diverging from Priors in RQ-2a

### v4

@@ Discussion
Applying our quality checks, we uncovered significant logical inconsistencies: specifically, most models (80%) produced “average” opinions more extreme than demographic-specific predictions, violating basic statistical logic. Unlike logical inconsistencies of LLMs exposed in previous research—for example, mutually incompatible factual beliefs (kassner2021beliefbank), self‑inconsistent chain‑of‑thought traces (wang2022selfconsistency), violations of basic propositional logic (ghosh2024logic), or within‑passage contradictions (mundler2024selfcontradictions)—these errors indicate fundamental logical flaws at the distributional level. Practitioners cannot simply acknowledge and proceed despite these errors; they must reconsider or even abandon flawed simulation strategies entirely.
Further, our checks reveal systematic discrepancies with stakeholder expectations. Older models (like GPT-3.5-turbo) tend to exaggerate opinion differences even on obviously non-divisive claims, risking harmful stereotyping or inefficient resource allocation. Conversely, newer models (GPT-4, GPT-4.1, GPT-5-mini, LLaMa3, Mistral, Titan) often obscure genuine demographic divides on controversial topics, potentially masking critical areas affecting particular groups. Importantly, we find that the magnitude of gender-level differences observed in QC2a correlates with the direction of the errors observed in QC2b, suggesting that testing clear-cut cases aligned with stakeholder priors—even in the absence of extensive human data—provides a valuable preliminary quality check of LLM reliability.
@@
Limitations
Our study provides a reusable framework for early‐stage evaluation of LLM‐simulated opinions, but several factors define the boundaries of its current scope and suggest directions for future work. First, in our empirical results we benchmark exclusively instruction–tuned language models, as prior literature suggests that such models are likely to perform better than others on instruction–style prompts. The proposed tests are agnostic to model type, but we leave evaluation of base models to future work. Second, our experiments use only two binary gender personas and an unprimed “average” persona as proof of concept. We constrain ourselves to binary gender due to limitations of the human data, and in Appendix H we show how QC1 can be extended to non-binary gender. Crucially, in many application domains the relevant demographic or psychographic spectrum is much richer. Identifying all important subgroups and capturing their diversity remains an open challenge; omitting a salient group can invalidate the average–persona check and obscure meaningful variation. Finally, the human annotator pool for our TopicMisinfo dataset is predominantly male ($\approx 69\,\%$ men, $31\,\%$ women). Although QC2b uses only subgroup–specific means and is therefore unaffected by the imbalance, this gender distribution could influence population‐level statistics if the dataset is reused for other purposes.
@@
Organizations should prioritize testing whether model outputs align with clear common-sense or domain-specific expectations, paying close attention to discrepancies with their justified priors. For instance, an error of exaggeration of differences between groups on a non-divisive issue is likely a more serious error than an error of exaggeration on an issue known to be divisive. Moreover, while our current metrics focus on mean differences in opinion distributions, some stakeholders may prioritize other distributional properties—such as variance or skew of opinions (bui2025mixture; wang2025large). Our framework is flexible: the test statistics we propose can be adapted to target other distributional moments.
@@ Conclusion
Acknowledgments.
This work was supported in part by Good Systems, a research grand challenge at the University of Texas at Austin. S.F. acknowledges support from the Schmidt Sciences AI2050 Early Career Fellowship.
Appendix A: Literature Review Table
@@ Acknowledgments.
[…]
Appendix B: Data Collection Procedure - TopicMisinfo Dataset
Using Amazon SageMaker Ground Truth, we hired American workers and compensated them at a rate of \$15-20 per hour to annotate various claims. We asked them to annotate each claim according to different important axes of prioritization. Specifically, we asked for annotations of a claim’s along four dimensions: (1) “Prioritization”: need for prioritization in fact-checking (based on their own subjective criteria); (2) “General Public”: interest to the general public; (3) “Group Harm”: potential to harm certain demographic groups; and (4) “Perceived Falsehood”: perceived degree of veracity. Table Appendix B: Data Collection Procedure - TopicMisinfo Dataset shows the questions that annotators were asked. If workers identified a claim as harmful to a specific group(s), we asked them to identify which group(s) may be harmed. After workers completed the task, we asked them to fill out a demographic survey.
@@
Appendix C: Example Claims from Dataset
@@
Appendix D: Descriptive Analysis of Dataset
@@
Appendix E: Prompts for Experiments
@@
Appendix F: Pseudocode for Quality Check 1a.
@@
Appendix G: Pseudocode for Quality Check 1b.
@@
8:    Initialize abs_diffs $\leftarrow[\;]$
@@
20: Initialize bootstrap_results $\leftarrow[\;]$
@@
Appendix H: Replication of Quality Check 1 including Non-binary Gender
@@
Appendix I: Fine-tuning approach for benchmark evaluation
@@
Appendix J: In-context learning approach for benchmark evaluation
@@
Appendix K: LLM Positioning of Opinions
This chart complements the results from Experiment 3, as seen in Tables 1 and 2, by showing the positioning of LLM-generated opinions on a Likert scale for men and women.
@@
Appendix L: Salient Examples of Opinion Differences Diverging from Priors in RQ-2a